# A Framework for Objective-Driven Dynamical Stochastic Fields

### v1

@@ Appendix B Proofs in Section 3
$\left\|e^{{{\mathbf{G}}}t}\Delta{\hat{\bm{\varphi}}}\right\|$ $=Z\cdot\left\|e^{{{\mathbf{G}}}t}({\hat{\bm{\varphi}}}_{1}-{\hat{\bm{\varphi}}}_{2})\right\|=Z\cdot\left\|e^{{{\mathbf{G}}}t}{\hat{\bm{\varphi}}}_{1}-{\hat{\bar{\bm{\varphi}}}}+{\hat{\bar{\bm{\varphi}}}}-e^{{{\mathbf{G}}}t}{\hat{\bm{\varphi}}}_{2}\right\|$ (B.31)
$\leq Z\cdot\left\|e^{{{\mathbf{G}}}t}{\hat{\bm{\varphi}}}_{1}-{\hat{\bar{\bm{\varphi}}}}\right\|+Z\cdot\left\|{\hat{\bar{\bm{\varphi}}}}-e^{{{\mathbf{G}}}t}{\hat{\bm{\varphi}}}_{2}\right\|$ (B.32)
@@
$\left\|{{\mathbf{S}}}\Delta{\hat{\bm{\varphi}}}\right\|$ $=\left\|\int_{0}^{\infty}{\mathop{}\!\mathrm{d}t}\ e^{{{\mathbf{G}}}t}\Delta{\hat{\bm{\varphi}}}\right\|\leq\int_{0}^{\infty}{\mathop{}\!\mathrm{d}t}\ \left\|e^{{{\mathbf{G}}}t}\Delta{\hat{\bm{\varphi}}}\right\|$ (B.34)
[…]
@@
Proposition 3.16 (Gradient Formula). Given an objective operator ${\bm{\Gamma}}:{\widetilde{\mathcal{H}}}\to{\mathbb{R}}$, assuming the system is ergodic, with the parameterization given by Definition 3.15 we have
$\frac{\partial{\bar{\gamma}}}{\partial G^{{\omega}^{\prime}}_{{\omega}}}={\bm{\Gamma}}(1+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}}){{\widetilde{\mathbf{A}}}}_{{\omega}^{\prime}}^{{\omega}}{\hat{\bar{\bm{\varphi}}}}.$ (B.36)
@@
$\delta{{\mathbf{G}}}=\epsilon\cdot{\mathbf{\Pi}}\ {{\widetilde{\mathbf{A}}}}_{{\omega}^{\prime}}^{{\omega}}H^{{\omega}^{\prime}}_{{\omega}}.$ (B.37)
@@
$\frac{\mathop{}\!\mathrm{d}}{{\mathop{}\!\mathrm{d}t}}{\hat{\bm{\varphi}}}=({{\mathbf{G}}}+\epsilon{{\mathbf{H}}}){\hat{\bm{\varphi}}},$ (B.40)
@@
${\hat{\bm{\varphi}}}(t)=e^{t({{\mathbf{G}}}+\epsilon{{\mathbf{H}}})}{\hat{\bm{\varphi}}}(0).$ (B.41)
@@
${\hat{\bm{\varphi}}}(t)=e^{t{{\mathbf{G}}}}{\hat{\bm{\varphi}}}(0)+\epsilon\int_{0}^{t}{\mathop{}\!\mathrm{d}s}\ e^{(t-s){{\mathbf{G}}}}{{\mathbf{H}}}{\hat{\bm{\varphi}}}(s).$ (B.42)
@@
${\hat{\bm{\varphi}}}(t)=\left(e^{t{{\mathbf{G}}}}+\epsilon\int_{0}^{t}{\mathop{}\!\mathrm{d}s}\ e^{(t-s){{\mathbf{G}}}}{{\mathbf{H}}}e^{s{{\mathbf{G}}}}\right){\hat{\bm{\varphi}}}(0)+{\mathcal{O}}(\epsilon^{2}).$ (B.43)
Since first order differential equation has a unique solution and that the initial point ${\hat{\bm{\varphi}}}(0)$ is arbitrary, it must be that
@@
We want to make sure that $\delta{\hat{\bar{\bm{\varphi}}}}={\mathcal{O}}(\epsilon)$ given the perturbation $\epsilon{\mathcal{H}}$ on the generator ${{\mathbf{G}}}$. This is indeed the case as we can see from the following argument. Given the ergodicity, ${\hat{\bar{\bm{\varphi}}}}$ is the unique solution to $(1+{{\mathbf{G}}}){\hat{\bm{\varphi}}}={\hat{\bm{\varphi}}}$, i.e., ${\hat{\bar{\bm{\varphi}}}}$ is the eigenvector correspond to the non-degenerate eigenvalue $1$ of the linear operator $(1+{{\mathbf{G}}})$. It is known that analytic perturbation of a linear operator leads to analytic perturbation of the eigenvectors corresponding to non-degenerate eigenvalues [Kato, 1995]. Therefore, the perturbation of $\delta{{\mathbf{G}}}=\epsilon{{\mathbf{H}}}$ leads to analytic perturbation of $\delta{\hat{\bar{\bm{\varphi}}}}={\mathcal{O}}(\epsilon)$.
@@
${\bar{\gamma}}={\bm{\Gamma}}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}.$ (B.49)
@@
$\delta{\bar{\gamma}}={\bm{\Gamma}}(\delta{\widetilde{\mathbf{G}}}){\hat{\bar{\bm{\varphi}}}}+{\bm{\Gamma}}{\widetilde{\mathbf{G}}}(\delta{\hat{\bar{\bm{\varphi}}}})+o(\epsilon),$ (B.50)
where $\delta{\widetilde{\mathbf{G}}}=\epsilon\cdot{{\widetilde{\mathbf{A}}}}_{{\omega}^{\prime}}^{{\omega}}H^{{\omega}^{\prime}}_{{\omega}}={\mathcal{O}}(\epsilon)$ and $\delta{\hat{\bar{\bm{\varphi}}}}={\mathcal{O}}(\epsilon)$. Then, by Lemma B.2,
$\delta{\hat{\bar{\bm{\varphi}}}}$ $=\lim_{t\to\infty}e^{({{\mathbf{G}}}+\delta{{\mathbf{G}}})t}{\hat{\bm{\varphi}}}(0)-e^{{{\mathbf{G}}}t}{\hat{\bm{\varphi}}}(0)=\lim_{t\to\infty}\left(e^{({{\mathbf{G}}}+\delta{{\mathbf{G}}})t}-e^{{{\mathbf{G}}}t}\right){\hat{\bm{\varphi}}}(0)$ (B.51)
$=\lim_{t\to\infty}\epsilon\cdot\int_{0}^{t}{\mathop{}\!\mathrm{d}s}\ e^{s{{\mathbf{G}}}}{{\mathbf{H}}}e^{(t-s){{\mathbf{G}}}}{\hat{\bm{\varphi}}}(0)+o(\epsilon).$ (B.52)
Given the ergodicity, the above holds for any initial state ${\hat{\bm{\varphi}}}(0)$. Thus, let us take ${\hat{\bm{\varphi}}}(0)={\hat{\bar{\bm{\varphi}}}}$. Noticing that ${\hat{\bar{\bm{\varphi}}}}$ is a fixed-point with respect to the system dynamics, i.e., $e^{(t-s){{\mathbf{G}}}}{\hat{\bar{\bm{\varphi}}}}={\hat{\bar{\bm{\varphi}}}}$, we have
$\delta{\hat{\bar{\bm{\varphi}}}}$ $=\lim_{t\to\infty}\epsilon\cdot\int_{0}^{t}{\mathop{}\!\mathrm{d}s}\ e^{s{{\mathbf{G}}}}{{\mathbf{H}}}e^{(t-s){{\mathbf{G}}}}{\hat{\bar{\bm{\varphi}}}}+o(\epsilon)$ (B.53)
$=\epsilon\cdot\int_{0}^{\infty}{\mathop{}\!\mathrm{d}s}\ e^{s{{\mathbf{G}}}}{{\mathbf{H}}}{\hat{\bar{\bm{\varphi}}}}+o(\epsilon)$ (B.54)
$=\epsilon\cdot{{\mathbf{S}}}{{\mathbf{H}}}{\hat{\bar{\bm{\varphi}}}}+o(\epsilon)$ (B.55)
$={{\mathbf{S}}}\delta{{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}+o(\epsilon).$ (B.56)
One can verify that the above is valid as ${{\mathbf{H}}}{\hat{\bar{\bm{\varphi}}}}={\mathbf{\Pi}}{{\widetilde{\mathbf{A}}}}_{{\omega}^{\prime}}^{{\omega}}H^{{\omega}^{\prime}}_{{\omega}}{\hat{\bar{\bm{\varphi}}}}\in{\mathbf{\Pi}}{\widetilde{\mathcal{H}}}$ indeed lies in ${\Delta\mathcal{H}}={\mathbf{\Pi}}{\widetilde{\mathcal{H}}}$, and thus ${{\mathbf{S}}}{{\mathbf{H}}}$ is bounded.
@@
$\delta{\bar{\gamma}}$ $={\bm{\Gamma}}\delta{{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}+{\bm{\Gamma}}{\widetilde{\mathbf{G}}}{{\mathbf{S}}}\delta{{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}+o(\epsilon)$ (B.57)
$=\epsilon\cdot{\bm{\Gamma}}{{\widetilde{\mathbf{A}}}}_{{\omega}^{\prime}}^{{\omega}}H^{{\omega}^{\prime}}_{{\omega}}{\hat{\bar{\bm{\varphi}}}}+\epsilon\cdot{\bm{\Gamma}}{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}}{{\widetilde{\mathbf{A}}}}_{{\omega}^{\prime}}^{{\omega}}H^{{\omega}^{\prime}}_{{\omega}}{\hat{\bar{\bm{\varphi}}}}+o(\epsilon)$ (B.58)
$=\epsilon\cdot{\bm{\Gamma}}(1+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}}){{\widetilde{\mathbf{A}}}}_{{\omega}^{\prime}}^{{\omega}}H^{{\omega}^{\prime}}_{{\omega}}{\hat{\bar{\bm{\varphi}}}}+o(\epsilon).$ (B.59)
@@
$\frac{\partial{\bar{\gamma}}}{\partial G^{{\omega}^{\prime}}_{{\omega}}}={\bm{\Gamma}}(1+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}}){{\widetilde{\mathbf{A}}}}_{{\omega}^{\prime}}^{{\omega}}{\hat{\bar{\bm{\varphi}}}}.$ (B.60)
@@
${\mathbf{\Pi}}\ {\widetilde{\mathbf{M}}}\ {\hat{\bm{e}}}_{\alpha\beta\mu\nu}$ $={\mathbf{\Pi}}\ M^{\alpha^{\prime}\mu^{\prime}}_{\alpha\beta\mu}{\hat{\bm{e}}}_{{\alpha^{\prime}\beta\mu^{\prime}\nu},{\alpha\beta\mu\nu}}$ (B.62)
$=\sum_{\alpha^{\prime}\mu^{\prime}}M^{\alpha^{\prime}\mu^{\prime}}_{\alpha\beta\mu}({\hat{\bm{e}}}_{{\alpha^{\prime}\beta\mu^{\prime}\nu}}-{\hat{\bm{e}}}_{\alpha\beta\mu\nu})$ (B.63)
$=\sum_{\alpha^{\prime}\mu^{\prime}:\alpha^{\prime}\mu^{\prime}\neq\alpha\mu}M^{\alpha^{\prime}\mu^{\prime}}_{\alpha\beta\mu}({\hat{\bm{e}}}_{{\alpha^{\prime}\beta\mu^{\prime}\nu}}-{\hat{\bm{e}}}_{\alpha\beta\mu\nu})$ (B.64)
$=\sum_{\alpha^{\prime}\mu^{\prime}}M^{\alpha^{\prime}\mu^{\prime}}_{\alpha\beta\mu}{\hat{\bm{e}}}_{{\alpha^{\prime}\beta\mu^{\prime}\nu}}$ (because $\sum_{\alpha^{\prime}\mu^{\prime}}M^{\alpha^{\prime}\mu^{\prime}}_{\alpha\beta\mu}=-M^{\alpha\mu}_{\alpha\beta\mu}$)
$={{\mathbf{M}}}{\hat{\bm{e}}}_{\alpha\beta\mu\nu}.$ (B.65)
@@
Proposition 3.19 (Gradient Formula in the Two-entity View). Given an objective operator ${\bm{\Gamma}}:{\widetilde{\mathcal{H}}}\to{\mathbb{R}}$, assuming the system is ergodic, with the parameterization of generator ${{\mathbf{M}}}$ given by Definition 3.17 we have
$\frac{\partial{\bar{\gamma}}}{\partial M^{\alpha^{\prime}\mu^{\prime}}_{\alpha\beta\mu}}={\bm{\Gamma}}(1+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}}){{\widetilde{\mathbf{A}}}}_{\alpha^{\prime}\mu^{\prime}}^{\alpha\beta\mu}{\hat{\bar{\bm{\varphi}}}}.$ (B.66)
@@
$\delta{\bar{\gamma}}$ $=\delta({\bm{\Gamma}}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}})={\bm{\Gamma}}(\delta{\widetilde{\mathbf{G}}}){\hat{\bar{\bm{\varphi}}}}+{\bm{\Gamma}}{\widetilde{\mathbf{G}}}(\delta{\hat{\bar{\bm{\varphi}}}})+o(\epsilon)$ (B.67)
$={\bm{\Gamma}}(\delta{\widetilde{\mathbf{G}}}){\hat{\bar{\bm{\varphi}}}}+{\bm{\Gamma}}{\widetilde{\mathbf{G}}}{{\mathbf{S}}}(\delta{{\mathbf{G}}}){\hat{\bar{\bm{\varphi}}}}+o(\epsilon).$ (B.68)
@@
$\delta{\widetilde{\mathbf{G}}}=\delta{\widetilde{\mathbf{M}}}=\delta M^{\alpha^{\prime}\mu^{\prime}}_{\alpha\beta\mu}{{\widetilde{\mathbf{A}}}}_{\alpha^{\prime}\mu^{\prime}}^{\alpha\beta\mu},\qquad\delta{{\mathbf{G}}}=\delta{{\mathbf{M}}}={\mathbf{\Pi}}\ \delta{\widetilde{\mathbf{M}}}={\mathbf{\Pi}}\ \delta M^{\alpha^{\prime}\mu^{\prime}}_{\alpha\beta\mu}{{\widetilde{\mathbf{A}}}}_{\alpha^{\prime}\mu^{\prime}}^{\alpha\beta\mu}.$ (B.69)
@@
$\delta{\bar{\gamma}}$ $={\bm{\Gamma}}(\delta M^{\alpha^{\prime}\mu^{\prime}}_{\alpha\beta\mu}{{\widetilde{\mathbf{A}}}}_{\alpha^{\prime}\mu^{\prime}}^{\alpha\beta\mu}){\hat{\bar{\bm{\varphi}}}}+{\bm{\Gamma}}{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}}\ (\delta M^{\alpha^{\prime}\mu^{\prime}}_{\alpha\beta\mu}{{\widetilde{\mathbf{A}}}}_{\alpha^{\prime}\mu^{\prime}}^{\alpha\beta\mu}){\hat{\bar{\bm{\varphi}}}}+o(\epsilon)$ (B.70)
$={\bm{\Gamma}}(1+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}})(\delta M^{\alpha^{\prime}\mu^{\prime}}_{\alpha\beta\mu}{{\widetilde{\mathbf{A}}}}_{\alpha^{\prime}\mu^{\prime}}^{\alpha\beta\mu}){\hat{\bar{\bm{\varphi}}}}+o(\epsilon).$ (B.71)
@@
$\frac{\partial{\bar{\gamma}}}{\partial M^{\alpha^{\prime}\mu^{\prime}}_{\alpha\beta\mu}}={\bm{\Gamma}}(1+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}}){{\widetilde{\mathbf{A}}}}_{\alpha^{\prime}\mu^{\prime}}^{\alpha\beta\mu}{\hat{\bar{\bm{\varphi}}}}.$ (B.72)

### v2

@@ Appendix B Proofs in Section 3
$\left\|e^{{{\mathbf{G}}}t}\Delta{\lvert\varphi\rangle}\right\|$ $=Z\cdot\left\|e^{{{\mathbf{G}}}t}({\lvert\varphi\rangle}_{1}-{\lvert\varphi\rangle}_{2})\right\|=Z\cdot\left\|e^{{{\mathbf{G}}}t}{\lvert\varphi\rangle}_{1}-{\lvert\bar{\varphi}\rangle}+{\lvert\bar{\varphi}\rangle}-e^{{{\mathbf{G}}}t}{\lvert\varphi\rangle}_{2}\right\|$ (B.31)
$\leq Z\cdot\left\|e^{{{\mathbf{G}}}t}{\lvert\varphi\rangle}_{1}-{\lvert\bar{\varphi}\rangle}\right\|+Z\cdot\left\|{\lvert\bar{\varphi}\rangle}-e^{{{\mathbf{G}}}t}{\lvert\varphi\rangle}_{2}\right\|$ (B.32)
@@
$\left\|{{\mathbf{S}}}\Delta{\lvert\varphi\rangle}\right\|$ $=\left\|\int_{0}^{\infty}{\mathop{}\!\mathrm{d}t}\ e^{{{\mathbf{G}}}t}\Delta{\lvert\varphi\rangle}\right\|\leq\int_{0}^{\infty}{\mathop{}\!\mathrm{d}t}\ \left\|e^{{{\mathbf{G}}}t}\Delta{\lvert\varphi\rangle}\right\|$ (B.34)
$\leq\int_{0}^{\infty}{\mathop{}\!\mathrm{d}t}\ 2ZKe^{-\kappa t}<\infty.$ (B.35)
@@
Proposition 3.16 (Gradient Formula). Given an objective operator $\bm{\Gamma}$$\widetilde{\mathcal{H}}$$\to{\mathbb{R}}$, assuming the system is ergodic, with the parameterization given by Definition 3.15 we have
$\frac{\partial{\bar{\gamma}}}{\partial G^{{\omega}^{\prime}}_{{\omega}}}={\bm{\Gamma}}(1+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}}){{\widetilde{\mathbf{A}}}}_{{\omega}^{\prime}}^{{\omega}}{\lvert\bar{\varphi}\rangle}.$ (B.36)
@@
$\delta{{\mathbf{G}}}=\epsilon\cdot\sum_{{\omega}{\omega}^{\prime}}{\mathbf{\Pi}}\ {{\widetilde{\mathbf{A}}}}_{{\omega}^{\prime}}^{{\omega}}H^{{\omega}^{\prime}}_{{\omega}}.$ (B.37)
@@
$\frac{\mathop{}\!\mathrm{d}}{{\mathop{}\!\mathrm{d}t}}{\lvert\varphi\rangle}=({{\mathbf{G}}}+\epsilon{{\mathbf{H}}}){\lvert\varphi\rangle},$ (B.40)
@@
$\lvert\varphi(t)\rangle=e^{t({{\mathbf{G}}}+\epsilon{{\mathbf{H}}})}\lvert\varphi(0)\rangle.$ (B.41)
@@
$\lvert\varphi(t)\rangle=e^{t{{\mathbf{G}}}}\lvert\varphi(0)\rangle+\epsilon\int_{0}^{t}{\mathop{}\!\mathrm{d}s}\ e^{(t-s){{\mathbf{G}}}}{{\mathbf{H}}}\lvert\varphi(s)\rangle.$ (B.42)
@@
$\lvert\varphi(t)\rangle=\left(e^{t{{\mathbf{G}}}}+\epsilon\int_{0}^{t}{\mathop{}\!\mathrm{d}s}\ e^{(t-s){{\mathbf{G}}}}{{\mathbf{H}}}e^{s{{\mathbf{G}}}}\right)\lvert\varphi(0)\rangle+{\mathcal{O}}(\epsilon^{2}).$ (B.43)
Since first order differential equation has a unique solution and that the initial point $\lvert\varphi(0)\rangle$ is arbitrary, it must be that
@@
We want to make sure that $\delta{\lvert\bar{\varphi}\rangle}={\mathcal{O}}(\epsilon)$ given the perturbation $\epsilon{\mathcal{H}}$ on the generator $\mathbf{G}$}}. This is indeed the case as we can see from the following argument. Given the ergodicity, $\lvert\bar{\varphi}\rangle$} is the unique solution to $(1+{{\mathbf{G}}}){\lvert\varphi\rangle}={\lvert\varphi\rangle}$, i.e., $\lvert\bar{\varphi}\rangle$} is the eigenvector correspond to the non-degenerate eigenvalue $1$ of the linear operator $(1+{{\mathbf{G}}})$. It is known that analytic perturbation of a linear operator leads to analytic perturbation of the eigenvectors corresponding to non-degenerate eigenvalues [Kato, 1995]. Therefore, the perturbation of $\delta{{\mathbf{G}}}=\epsilon{{\mathbf{H}}}$ leads to analytic perturbation of $\delta{\lvert\bar{\varphi}\rangle}={\mathcal{O}}(\epsilon)$.
@@
${\bar{\gamma}}={\bm{\Gamma}}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}.$ (B.49)
@@
$\delta{\bar{\gamma}}={\bm{\Gamma}}(\delta{\widetilde{\mathbf{G}}}){\lvert\bar{\varphi}\rangle}+{\bm{\Gamma}}{\widetilde{\mathbf{G}}}(\delta{\lvert\bar{\varphi}\rangle})+o(\epsilon),$ (B.50)
where $\delta{\widetilde{\mathbf{G}}}=\epsilon\cdot\sum_{{\omega}{\omega}^{\prime}}{{\widetilde{\mathbf{A}}}}_{{\omega}^{\prime}}^{{\omega}}H^{{\omega}^{\prime}}_{{\omega}}={\mathcal{O}}(\epsilon)$ and $\delta{\lvert\bar{\varphi}\rangle}={\mathcal{O}}(\epsilon)$. Then, by Lemma B.2,
$\delta{\lvert\bar{\varphi}\rangle}$ $=\lim_{t\to\infty}e^{({{\mathbf{G}}}+\delta{{\mathbf{G}}})t}\lvert\varphi(0)\rangle-e^{{{\mathbf{G}}}t}\lvert\varphi(0)\rangle=\lim_{t\to\infty}\left(e^{({{\mathbf{G}}}+\delta{{\mathbf{G}}})t}-e^{{{\mathbf{G}}}t}\right)\lvert\varphi(0)\rangle$ (B.51)
$=\lim_{t\to\infty}\epsilon\cdot\int_{0}^{t}{\mathop{}\!\mathrm{d}s}\ e^{s{{\mathbf{G}}}}{{\mathbf{H}}}e^{(t-s){{\mathbf{G}}}}\lvert\varphi(0)\rangle+o(\epsilon).$ (B.52)
Given the ergodicity, the above holds for any initial state $\lvert\varphi(0)\rangle$. Thus, let us take $\lvert\varphi(0)\rangle={\lvert\bar{\varphi}\rangle}$. Noticing that $\lvert\bar{\varphi}\rangle$} is a fixed-point with respect to the system dynamics, i.e., $e^{(t-s){{\mathbf{G}}}}{\lvert\bar{\varphi}\rangle}={\lvert\bar{\varphi}\rangle}$, we have
$\delta{\lvert\bar{\varphi}\rangle}$ $=\lim_{t\to\infty}\epsilon\cdot\int_{0}^{t}{\mathop{}\!\mathrm{d}s}\ e^{s{{\mathbf{G}}}}{{\mathbf{H}}}e^{(t-s){{\mathbf{G}}}}{\lvert\bar{\varphi}\rangle}+o(\epsilon)$ (B.53)
$=\epsilon\cdot\int_{0}^{\infty}{\mathop{}\!\mathrm{d}s}\ e^{s{{\mathbf{G}}}}{{\mathbf{H}}}{\lvert\bar{\varphi}\rangle}+o(\epsilon)$ (B.54)
$=\epsilon\cdot{{\mathbf{S}}}{{\mathbf{H}}}{\lvert\bar{\varphi}\rangle}+o(\epsilon)$ (B.55)
$={{\mathbf{S}}}\delta{{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}+o(\epsilon).$ (B.56)
One can verify that the above is valid as ${{\mathbf{H}}}{\lvert\bar{\varphi}\rangle}={\mathbf{\Pi}}\sum_{{\omega}{\omega}^{\prime}}{{\widetilde{\mathbf{A}}}}_{{\omega}^{\prime}}^{{\omega}}H^{{\omega}^{\prime}}_{{\omega}}{\lvert\bar{\varphi}\rangle}\in{\mathbf{\Pi}}{\widetilde{\mathcal{H}}}$ indeed lies in ${\Delta\mathcal{H}}={\mathbf{\Pi}}{\widetilde{\mathcal{H}}}$, and thus ${{\mathbf{S}}}{{\mathbf{H}}}$ is bounded.
@@
$\delta{\bar{\gamma}}$ $={\bm{\Gamma}}\delta{{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}+{\bm{\Gamma}}{\widetilde{\mathbf{G}}}{{\mathbf{S}}}\delta{{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}+o(\epsilon)$ (B.57)
$=\epsilon\cdot{\bm{\Gamma}}\sum_{{\omega}{\omega}^{\prime}}{{\widetilde{\mathbf{A}}}}_{{\omega}^{\prime}}^{{\omega}}H^{{\omega}^{\prime}}_{{\omega}}{\lvert\bar{\varphi}\rangle}+\epsilon\cdot{\bm{\Gamma}}{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}}\sum_{{\omega}{\omega}^{\prime}}{{\widetilde{\mathbf{A}}}}_{{\omega}^{\prime}}^{{\omega}}H^{{\omega}^{\prime}}_{{\omega}}{\lvert\bar{\varphi}\rangle}+o(\epsilon)$ (B.58)
$=\epsilon\cdot{\bm{\Gamma}}(1+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}})\sum_{{\omega}{\omega}^{\prime}}{{\widetilde{\mathbf{A}}}}_{{\omega}^{\prime}}^{{\omega}}H^{{\omega}^{\prime}}_{{\omega}}{\lvert\bar{\varphi}\rangle}+o(\epsilon).$ (B.59)
@@
$\frac{\partial{\bar{\gamma}}}{\partial G^{{\omega}^{\prime}}_{{\omega}}}={\bm{\Gamma}}(1+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}}){{\widetilde{\mathbf{A}}}}_{{\omega}^{\prime}}^{{\omega}}{\lvert\bar{\varphi}\rangle}.$ (B.60)
@@
${\mathbf{\Pi}}\ {\widetilde{\mathbf{M}}}\ \lvert\alpha\beta\mu\nu\rangle$ $={\mathbf{\Pi}}\ M^{\alpha^{\prime}\mu^{\prime}}_{\alpha\beta\mu}\lvert{\alpha^{\prime}\beta\mu^{\prime}\nu},{\alpha\beta\mu\nu}\rangle$ (B.62)
$=\sum_{\alpha^{\prime}\mu^{\prime}}M^{\alpha^{\prime}\mu^{\prime}}_{\alpha\beta\mu}(\lvert{\alpha^{\prime}\beta\mu^{\prime}\nu}\rangle-\lvert\alpha\beta\mu\nu\rangle)$ (B.63)
$=\sum_{\alpha^{\prime}\mu^{\prime}:\alpha^{\prime}\mu^{\prime}\neq\alpha\mu}M^{\alpha^{\prime}\mu^{\prime}}_{\alpha\beta\mu}(\lvert{\alpha^{\prime}\beta\mu^{\prime}\nu}\rangle-\lvert\alpha\beta\mu\nu\rangle)$ (B.64)
$=\sum_{\alpha^{\prime}\mu^{\prime}}M^{\alpha^{\prime}\mu^{\prime}}_{\alpha\beta\mu}\lvert{\alpha^{\prime}\beta\mu^{\prime}\nu}\rangle$ (because $\sum_{\alpha^{\prime}\mu^{\prime}}M^{\alpha^{\prime}\mu^{\prime}}_{\alpha\beta\mu}=-M^{\alpha\mu}_{\alpha\beta\mu}$)
$={{\mathbf{M}}}\lvert\alpha\beta\mu\nu\rangle.$ (B.65)
@@
Proposition 3.19 (Gradient Formula in the Two-entity View). Given an objective operator $\bm{\Gamma}$$\widetilde{\mathcal{H}}$$\to{\mathbb{R}}$, assuming the system is ergodic, with the parameterization of generator $\mathbf{M}$}} given by Definition 3.17 we have
$\frac{\partial{\bar{\gamma}}}{\partial M^{\alpha^{\prime}\mu^{\prime}}_{\alpha\beta\mu}}={\bm{\Gamma}}(1+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}}){{\widetilde{\mathbf{A}}}}_{\alpha^{\prime}\mu^{\prime}}^{\alpha\beta\mu}{\lvert\bar{\varphi}\rangle}.$ (B.66)
@@
$\delta{\bar{\gamma}}$ $=\delta({\bm{\Gamma}}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle})={\bm{\Gamma}}(\delta{\widetilde{\mathbf{G}}}){\lvert\bar{\varphi}\rangle}+{\bm{\Gamma}}{\widetilde{\mathbf{G}}}(\delta{\lvert\bar{\varphi}\rangle})+o(\epsilon)$ (B.67)
$={\bm{\Gamma}}(\delta{\widetilde{\mathbf{G}}}){\lvert\bar{\varphi}\rangle}+{\bm{\Gamma}}{\widetilde{\mathbf{G}}}{{\mathbf{S}}}(\delta{{\mathbf{G}}}){\lvert\bar{\varphi}\rangle}+o(\epsilon).$ (B.68)
@@
$\delta{\widetilde{\mathbf{G}}}=\delta{\widetilde{\mathbf{M}}}=\sum_{\alpha^{\prime}\beta^{\prime}\alpha\beta\mu}\delta M^{\alpha^{\prime}\mu^{\prime}}_{\alpha\beta\mu}{{\widetilde{\mathbf{A}}}}_{\alpha^{\prime}\mu^{\prime}}^{\alpha\beta\mu},\qquad\delta{{\mathbf{G}}}=\delta{{\mathbf{M}}}={\mathbf{\Pi}}\ \delta{\widetilde{\mathbf{M}}}={\mathbf{\Pi}}\ \sum_{\alpha^{\prime}\beta^{\prime}\alpha\beta\mu}\delta M^{\alpha^{\prime}\mu^{\prime}}_{\alpha\beta\mu}{{\widetilde{\mathbf{A}}}}_{\alpha^{\prime}\mu^{\prime}}^{\alpha\beta\mu}.$ (B.69)
@@
$\delta{\bar{\gamma}}$ $={\bm{\Gamma}}(\sum_{\alpha^{\prime}\beta^{\prime}\alpha\beta\mu}\delta M^{\alpha^{\prime}\mu^{\prime}}_{\alpha\beta\mu}{{\widetilde{\mathbf{A}}}}_{\alpha^{\prime}\mu^{\prime}}^{\alpha\beta\mu}){\lvert\bar{\varphi}\rangle}+{\bm{\Gamma}}{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}}\ (\sum_{\alpha^{\prime}\beta^{\prime}\alpha\beta\mu}\delta M^{\alpha^{\prime}\mu^{\prime}}_{\alpha\beta\mu}{{\widetilde{\mathbf{A}}}}_{\alpha^{\prime}\mu^{\prime}}^{\alpha\beta\mu}){\lvert\bar{\varphi}\rangle}+o(\epsilon)$ (B.70)
$={\bm{\Gamma}}(1+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}})(\sum_{\alpha^{\prime}\beta^{\prime}\alpha\beta\mu}\delta M^{\alpha^{\prime}\mu^{\prime}}_{\alpha\beta\mu}{{\widetilde{\mathbf{A}}}}_{\alpha^{\prime}\mu^{\prime}}^{\alpha\beta\mu}){\lvert\bar{\varphi}\rangle}+o(\epsilon).$ (B.71)
@@
$\frac{\partial{\bar{\gamma}}}{\partial M^{\alpha^{\prime}\mu^{\prime}}_{\alpha\beta\mu}}={\bm{\Gamma}}(1+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}}){{\widetilde{\mathbf{A}}}}_{\alpha^{\prime}\mu^{\prime}}^{\alpha\beta\mu}{\lvert\bar{\varphi}\rangle}.$ (B.72)